If you are given 100feet of fence, what dimensions would give you the maximum area of a rectangular fenced in garden plot?
Found 2 solutions by checkley77, JBarnum:
Answer by checkley77(12844)

The maximum area for a rectangle is a SQUARE.
Thus for 100 ft. of fencing each side would be 100/4=25 ft.
25^2=625 ft^2 ans.
Proof:
Let the sides be 24 & 26 ft.
24*26=624 ft^2.

Answer by JBarnum(2146)

a rectangle L=26 and W=24 gives A=624 26+26+24+24=100
a square L=25 and W=25 gives A=625 25+25+25+25=100
if u count the square as a rectangle then its 25x25